# Pluricanonical Geometry of Varieties Isogenous to a Product: Chevalley-Weil Theory and Pluricanonical Decompositions of Abelian Covers

### v1

@@ Abstract.
with $g(C_{i})\geq 2$ and $G$ acting freely. We establish the Chevalley–Weil formula for pluricanonical representations of a curve with a finite group action and a decomposition theorem for pluricanonical systems of abelian covers. These tools allow an explicit study of geometric properties of $X$, such as base loci and the birationality of pluricanonical maps. For threefolds isogenous to a product, we prove that the 4-canonical map is birational for $p_{g}\geq 5$ and construct an example attaining the maximal canonical degree for this class of threefolds. In this example, the canonical map is the normalization of its image, which admits isolated non-normal singularities. Computational classifications also reveal threefolds where the bicanonical map fails to be birational, even in the absence of genus-2 fibrations. This illustrates an interesting phenomenon similar to the non-standard case for surfaces.
@@
Acknowledgments. The authors would like to thank Roberto Pignatelli and Federico Fallucca for helpful comments and feedback. In particular, they are grateful to Roberto Pignatelli for helping with the proof of Theorem 5.6, and also for discussions concerning the content of Theorem 4.7. The first-named author would also like to thank Vladimir Lazić, Aryaman Patel and Francesco Antonio Denisi for engaging discussions and valuable insights.
@@
1. Introduction
Canonical and pluricanonical maps are fundamental objects in birational geometry and provide key information on the geometry of complex varieties of general type. It is classically known that for a smooth projective curve $C$ of genus $g(C)\geq 2$ the canonical system $|K_{C}|$ defines a morphism
$\Phi_{K_{C}}\colon C\longrightarrow\mathbb{P}_{\mathbb{C}}^{g(C)-1}$
called the canonical map. This map is an embedding if and only if $C$ is non-hyperelliptic; if $C$ is hyperelliptic, $\Phi_{K_{C}}$ is a finite morphism of degree two onto the rational normal curve.
@@ 1. Introduction
so that techniques from the theory of abelian covers [Par91], combined with representation-theoretic methods such as the Chevalley–Weil formula [CheWeil34], become available to analyze the canonical system in detail. In the same spirit, Catanese [Cat18] used this class of surfaces to construct examples with birational canonical maps and canonical images of high degree. This motivates our systematic study of canonical and pluricanonical maps of varieties isogenous to a product of unmixed type, i.e., smooth projective varieties of the form
@@
where each $C_{i}$ is a smooth projective curve of genus at least $2$ and $G$ is a finite group acting freely and diagonally on the product. These objects naturally generalize surfaces isogenous to a product of unmixed type to higher dimensions; for brevity, we will often refer to them as VIPs. Our main focus is to investigate the birationality of their associated canonical and pluricanonical maps. When attempting to extend to VIPs and their pluricanonical maps the above-mentioned strategy, which combines representation-theoretic methods and the theory of abelian covers, two significant challenges arise.
First, while the Chevalley–Weil formula [CheWeil34] is a classical tool for understanding the canonical representation of a Riemann surface $C$ with an action of a finite group $G$, a detailed treatment for pluricanonical representations does not seem to appear in the literature. Although the statement is mentioned, for example, in [Li18, Section 5], a proof is not given there. For this reason, we provide a precise formulation and a complete proof of such a formula.
@@
$\varphi_{m}\colon G\to\operatorname{GL}(H^{0}(C,K_{C}^{\otimes m}))$
and $\chi\in\operatorname{Irr}(G)$ be the character of an irreducible representation $\varrho\colon G\to\mathrm{GL}(V)$. If $m\geq 2$, then it holds
Here, $n_{1},\dotso,n_{r}$ are the branching indices of the cover $C\to C/G$, the integer $b]_{n_{i}}\in\{0,\ldots,n_{i}-1\}$ is the congruence class of $b$ modulo $n_{i}$ and $N_{i,\alpha}$ denotes the multiplicity of $\zeta_{n_{i}}^{\alpha}$ as an eigenvalue of $\varrho(h_{i})$, where $h_{1},\dotso,h_{r}$ are the local monodromies.
Second, we want to use Pardini’s theory of abelian covers to get a decomposition of the $m$-canonical system $H^{0}(X,K_{X}^{\otimes m})$ in a way that allows us to study geometric features, such as base points, rather than just the character of the Galois action. Such decompositions have been studied explicitly in the literature for $m=1$ by Pardini [Par91] and Liedtke [Lie03], and for $m=2$ by Bauer–Pignatelli [BP21]. Since we could not find a reference in the literature for $m\geq 3$, we provide a general version of such a decomposition formula for finite abelian covers and arbitrary $m$. This is the content of the following theorem. [^1] In a private communication, Roberto Pignatelli informed us that a similar version of Theorem B was presented by him in a talk at the University of Bayreuth in 2022. We point out that our proof is different and explicitly provides the correction terms $t_{(H,\psi)}^{r(m,H,\psi,\chi)}$. This explicit description is important for some of our computations, for instance to study the base loci of the pluricanonical systems.
Theorem B (Theorem 4.7).
@@
Theorem A and Theorem B provide us with the technical tools to analyze the pluricanonical maps and systems of an unmixed VIP with an abelian group in concrete examples.
After giving some generalities on the (pluri)canonical maps of VIPs, we specialize to the three-dimensional case. Here, we provide a general result on the birationality of the $4$-canonical map, which is based on the work of Meng Chen and Yong Hu [CH21] and on the geometry of Horikawa’s (1,2) surfaces [Hor76II]. According to [CCZ07], the $m$-canonical map of a minimal Gorenstein projective threefold of general type with canonical singularities is birational for $m\geq 5$.
Theorem C (Theorem 5.6).
Let $X$ be a threefold isogenous to a product of curves. If $p_{g}(X)\geq 5$, then the $4$-canonical map is birational.
Next, we focus on providing numerous explicit examples. In particular, we implement a classification algorithm for regular unmixed three-dimensional VIPs with an abelian group. Our approach combines modifications of the algorithms presented in [FG16] and [GPR22] with implementations of Theorem A, of Theorem B and of a generalization to higher dimensions and to pluricanonical maps of the representation-theoretic birationality criteria (see Proposition 5.13 and Corollary 5.14) originally introduced by Catanese in the surface case for the canonical map [Cat18]. More precisely, given an integer $\chi\leq-1$, our algorithm classifies all unmixed three-dimensional VIPs with holomorphic Euler-Poincaré characteristic $\chi$ and with abelian group $G$ such that the $G$-action on each curve is faithful. We run it for $\chi$ in the range $8\leq\chi\leq-1$. The output shows that no threefolds in the range $7\leq\chi\leq-1$ satisfy the above-mentioned birationality criteria for the canonical map. However, examples with a birational canonical map do exist for $\chi=-8$. We analyze in detail one such example, which leads to the following theorem.
Theorem D (Theorem 6.2).
@@
Finally, we run a classification algorithm which allows non-faithful actions of the group $G$ on the curves $C_{i}$. Applying it for $\chi=-1$ and $|G|\leq 16$, we obtain $347$ families of threefolds, summarized in Table LABEL:Table-chi-1. The table lists the Galois group $G$, the orders of the kernels, the branching types of the Galois covers $C_{i}\to C_{i}/G_{i}\simeq\mathbb{P}^{1}$ and the Hodge numbers of the threefolds. Additionally, for the canonical and bicanonical maps, it indicates the number of families for which these maps are base-point free, birational, non-birational and the number of those for which we cannot decide about the birationality by using our criteria. We also comment on the failure of birationality of the bicanonical map in our examples. To put things into perspective, we first recall the situation in lower dimensions. For a curve $C$, the bicanonical map is always an embedding unless $g(C)=2$; in this case, it is a double cover of a conic in $\mathbb{P}^{2}$. In dimension two, the works of Bombieri [Bomb73] and Reider [Reider88] imply that the bicanonical map of a minimal surface $S$ of general type with $K_{S}^{2}\geq 10$ is birational, except when $S$ admits a genus-2 fibration. This is the so-called standard case; see also the survey [BCP06, Section 2] for an overview on the topic. In our setting, the bicanonical map of some threefolds listed in Table LABEL:Table-chi-1 fails to be birational for the presence of a genus-2 fibration: this is similar to the standard case. However, we also identify and highlight examples without any genus-2 fibration whose bicanonical map is nevertheless non-birational, see Remark 6.3. This phenomenon is particularly interesting, as it is similar to the non-standard case for surfaces.
@@
We make one last observation. For all examples listed in Table LABEL:Table-chi-1, $K_{X}^{\otimes 2}$ is base-point free and $K_{X}^{\otimes 3}$ defines a birational map. This raises the natural questions of whether these properties hold for all regular unmixed three-dimensional VIPs.
Our paper is structured as follows.
In Section 2, we review the basic theory of varieties isogenous to a product of curves. A key ingredient is the existence and uniqueness of a minimal realization of a VIP. This allows us to describe these varieties in terms of certain combinatorial data associated with the group, the so-called algebraic datum. We focus on the unmixed case, where the group action is diagonal.
@@
In Section 3, we study the characters of the pluricanonical representations of a VIP and prove Theorem A.
@@
In Section 4, we specialize to the abelian case. Here, we prove Theorem B and then apply it to describe the pluricanonical systems of unmixed VIPs with an abelian group.
@@
In Section 5, we apply the theory developed in the preceding sections to study the birationality of pluricanonical maps of unmixed VIPs with an abelian group. We begin with general observations on the birationality of these maps, and then prove Theorem C. Finally, we discuss Catanese’s birationality criteria, adapting them to the context of pluricanonical maps and higher-dimensional varieties.
@@
In Section 6, we describe in detail the example leading to Theorem D. Finally, we present and discuss the above-mentioned computational results, see Table LABEL:Table-chi-1 and Remark 6.3.
2. Varieties Isogenous to a Product of Unmixed Type
@@ 2.1. Group-Theoretic Description of an Unmixed VIP
$\overline{\psi_{i}}\colon G_{i}\to\operatorname{Aut}(C_{i})$
@@
is the set of group elements of $H$ with non-trivial fixed locus on the curve $C$. It is hence called the stabilizer set of $V$.
@@ 3. Pluricanonical Representations of an Unmixed VIP
(1) ${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${C_{1}\times\dotso\times C_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{1}^{\prime}\times\dotso\times C_{n}^{\prime}}$
@@
(2) $(g_{1},1,\dotso,1)\cdot(g,g,\dotso,g)\cdot(g_{1}^{-1},1,\dotso,1)\cdot(h^{-1},h^{-1},\dotso,h^{-1})\in K_{1}\times\dotso\times K_{n}.$
In other words, by the minimality of the realization we have that $gh^{-1}\in K_{2}\cap\dotso\cap K_{n}=\{1\}.$ Thus, $g=h$, which implies by (2) that $g_{1},g]\in K_{1}$. Since $g_{1},g\in G$ are arbitrary, this shows that $G,G]\leq K_{1}$. Similarly, one can show that $G,G]\leq K_{i}$, $i=2,\dotso,n$. Thus, we get that $G,G]=\bigcap_{i=1}^{n}K_{i}=\{1\}$, i.e. $G$ is abelian. ∎
@@
(3) $\rho\colon G_{1}\times\dotso\times G_{n}\to\operatorname{GL}\big(H^{0}(C_{1}\times\dotso\times C_{n},K_{C_{1}\times\dotso\times C_{n}}^{\otimes m})\big).$
@@
As for (1), let $v\in V^{H}$ and $g\in N$ and $h\in H$. Then $g^{-1}hg\in H$, and therefore $\rho(g^{-1}hg)(v)=v$, which implies $\rho(h)\rho(g)(v)=\rho(g)(v)$, i.e. $\rho(g)(v)\in V^{H}$.
As for (2), by Maschke’s Lemma there exists an $N$-invariant complement $W$ of $V^{H}$, namely $V=V^{H}\oplus W$. A constituent $n_{\chi}\chi$ of $\chi_{\rho_{|N}}$ belongs to the character of $V^{H}$ if and only if $\chi_{|H}$ is a multiple of the trivial character $\chi_{triv}$.
By Künneth formula the action $\rho$ from (3) is the tensor product of the representations
@@
whose characters can be computed by the Chevalley-Weil formula in terms of the generating vectors $V_{i}$ of an algebraic datum of $X$. If $m=1$, we refer to the classical formula [CheWeil34] (see [FG16] for a modern proof), while for the case $m\geq 2$ we provide a generalization of the latter to pluricanonical representations in the subsection below. Thus, by the Chevalley-Weil formula and by Lemma 3.3 we recover from an algebraic datum of $X$ the character $\chi_{m}$ of the representation $\rho_{m}$.

### v2

# Pluricanonical Geometry of Varieties Isogenous to a Product and Pluricanonical Decompositions of Abelian Covers

@@ Abstract.
with $g(C_{i})\geq 2$ and $G$ acting freely. For this purpose, we provide a technical result which is of general interest: a decomposition theorem for pluricanonical systems of abelian covers. This theorem provides an effective tool for the explicit study of geometric properties, such as base loci and the birationality of pluricanonical maps.
For threefolds isogenous to a product, we prove that the 4-canonical map is birational for $p_{g}\geq 5$ and construct an example attaining the maximal canonical degree for this class of threefolds. In this example, the canonical map is the normalization of its image, which admits isolated non-normal singularities. Computational classifications also reveal threefolds where the bicanonical map fails to be birational, even in the absence of genus-2 fibrations. This illustrates an interesting phenomenon similar to the non-standard case for surfaces.
@@
Acknowledgments. The authors would like to thank Roberto Pignatelli and Federico Fallucca for helpful comments and feedback. In particular, they are grateful to Roberto Pignatelli for helping with the proof of Theorem 1.12, and also for discussions concerning the content of Theorem 3.7. They are also indebted to Vladimir Lazić for his help with the proof of Proposition 5.4.
The authors further thank Bernhard Köck and Wenfei Liu for useful comments and references on the Chevalley-Weil formula.
The first-named author would like to thank Aryaman Patel and Francesco Antonio Denisi for engaging discussions and valuable insights.
@@
Introduction
Canonical and pluricanonical maps are fundamental objects in birational geometry and provide key information on the geometry of complex varieties of general type. Recall that for a smooth projective curve $C$ of genus $g(C)\geq 2$ the canonical system $|K_{C}|$ defines a morphism
$\Phi_{K_{C}}\colon C\longrightarrow\mathbb{P}_{\mathbb{C}}^{g(C)-1}.$
This is an embedding if and only if $C$ is non-hyperelliptic; if $C$ is hyperelliptic, $\Phi_{K_{C}}$ is a finite morphism of degree two onto the rational normal curve.
@@ Introduction
so that techniques from the theory of abelian covers (see [Par91]), combined with representation-theoretic methods, become available to analyze the canonical system in detail. In the same spirit, Catanese [Cat18] used this class of surfaces to construct examples with birational canonical maps and canonical images of high degree. This motivates our systematic study of canonical and pluricanonical maps of varieties isogenous to a product of unmixed type, i.e., smooth projective varieties of the form
@@
where each $C_{i}$ is a smooth projective curve of genus at least $2$ and $G$ is a finite group acting freely and diagonally on the product. These objects naturally generalize surfaces isogenous to a product of unmixed type to higher dimensions; for brevity, we will often refer to them as VIPs.
Our main focus is to investigate the birationality of their associated canonical and pluricanonical maps. When attempting to extend to VIPs and their pluricanonical maps the above-mentioned strategy, significant challenges arise. Recall that the Chevalley-Weil formula [CheWeil34, Weil35] (see also [EL80], [K04], [K05], [LL25a], [LL25b] for various generalizations) allows one to determine the characters of the canonical and pluricanonical representations of a Riemann surface $C$ endowed with an action of a finite group $G$. While this information is useful for the study of the $m$-canonical linear system $H^{0}(X,K_{X}^{\otimes m})$, it is by itself insufficient to extract finer geometric properties, such as the presence of base points. To overcome this limitation in the case where $G$ is abelian, we use Pardini’s theory of abelian covers to obtain a complete and detailed decomposition of the pluricanonical systems, rather than just the character of the Galois action. Such decompositions have been studied explicitly in the literature for $m=1$ by Pardini [Par91] and Liedtke [Lie03], and for $m=2$ by Bauer–Pignatelli [BP21]. Since we could not find a reference in the literature for $m\geq 3$, we provide a general version of such a decomposition formula for finite abelian covers and arbitrary $m$, which we believe to be of independent interest. This is the content of the following theorem. [^1] In a private communication, Roberto Pignatelli informed us that a similar version of Theorem A was presented by him in a talk at the University of Bayreuth in 2022. We point out that our proof is different and explicitly provides the correction terms $t_{(H,\psi)}^{r(m,H,\psi,\chi)}$. This explicit description is important for some of our computations, for instance to study the base loci of the pluricanonical systems.
@@
Next, we give a generalization to higher dimensions and to pluricanonical maps of the representation-theoretic birationality criteria (see Proposition 4.7 and Corollary 4.8) originally introduced by Catanese [Cat18] in the surface case for the canonical map.
These criteria togheter with the Chevalley-Weil formula (see Theorem 2.4) and Theorem A provide us with the technical tools to analyze the pluricanonical maps and systems of an unmixed VIP with abelian Galois group in concrete examples. More precisely, we implement a classification algorithm for regular unmixed three-dimensional VIPs with abelian Galois group. Our approach combines modifications of the algorithms presented in [FG16] and [GPR22] with implementations of the above-mentioned tools: given an integer $\chi\leq-1$, our algorithm classifies all unmixed three-dimensional VIPs with holomorphic Euler-Poincaré characteristic $\chi$ and with abelian group $G$ such that the $G$-action on each curve is faithful. We run it for $\chi$ in the range $8\leq\chi\leq-1$. The output shows that no threefolds in the range $7\leq\chi\leq-1$ satisfy the above-mentioned birationality criteria for the canonical map. However, examples with a birational canonical map do exist for $\chi=-8$. We analyze in detail one such example, which leads to the following theorem.
Theorem B (Theorem 5.2).
@@
We also run a classification algorithm which allows non-faithful actions of the group $G$ on the curves $C_{i}$. Applying it for $\chi=-1$ and $|G|\leq 16$, we obtain $347$ families of threefolds, summarized in Table LABEL:Table-chi-1. The table lists the Galois group $G$, the orders of the kernels, the branching types of the Galois covers $C_{i}\to C_{i}/G_{i}\simeq\mathbb{P}^{1}$ and the Hodge numbers of the threefolds. Additionally, for the canonical and bicanonical maps, it indicates the number of families for which these maps are base-point free, birational, non-birational and the number of those for which we cannot decide about the birationality by using our criteria. We also comment on the failure of birationality of the bicanonical map in our examples. To put things into perspective, we recall the situation in lower dimensions. For a curve $C$, the bicanonical map is always an embedding unless $g(C)=2$; in this case, it is a double cover of a conic in $\mathbb{P}^{2}$. In dimension two, the works of Bombieri [Bomb73] and Reider [Reider88] imply that the bicanonical map of a minimal surface $S$ of general type with $K_{S}^{2}\geq 10$ is birational, except when $S$ admits a genus-2 fibration. This is the so-called standard case; see also the survey [BCP06, Section 2] for an overview on the topic. In our setting, the bicanonical map of some threefolds listed in Table LABEL:Table-chi-1 fails to be birational for the presence of a genus-2 fibration: this is similar to the standard case. However, we also identify and highlight examples without any genus-2 fibration whose bicanonical map is nevertheless non-birational, see Remark 5.3. This phenomenon is particularly interesting, as it is similar to the non-standard case for surfaces.
@@
We further observe that for all examples listed in Table LABEL:Table-chi-1, $K_{X}^{\otimes 2}$ is base-point free and $K_{X}^{\otimes 3}$ defines a birational map. This raises the natural questions of whether these properties hold for all regular unmixed three-dimensional VIPs.
Besides constructing and analyzing explicit examples, we also establish general results concerning the (pluri)canonical maps of varieties isogenous to a product, which are valid even when the group action is non-diagonal. In particular, we prove a birationality result for the $4$-canonical map of threefolds isogenous to a product. Our argument relies on the work of Chen-Hu [CH21] and on the geometry of Horikawa $(1,2)$-surfaces [Hor76II]. Recall that, by [CCZ07], the $m$-canonical map of a minimal Gorenstein projective threefold of general type with canonical singularities is birational for $m\geq 5$.
Theorem C (Theorem 1.12).
Let $X$ be a threefold isogenous to a product of curves. If $p_{g}(X)\geq 5$, then the $4$-canonical map of $X$ is birational.
The paper is structured as follows.
In Section 1, we review the basic theory of varieties isogenous to a product of curves. We also establish general results concerning the pluricanonical maps of VIPs and prove Theorem C.
@@
In Section 2, we study the characters of the pluricanonical representations of a VIP and give a proof of the Chevalley-Weil formula for $m$-canonical representations with $m\geq 2$ (see Theorem 2.4).
@@
In Section 3, we focus on abelian covers. Here, we prove Theorem A and then apply it to describe the pluricanonical systems of unmixed VIPs with abelian Galois group.
@@
In Section 4, we discuss Catanese’s birationality criteria, adapting them to the context of pluricanonical maps and higher-dimensional varieties.
@@
In Section 5, we describe in detail the example leading to Theorem B. Finally, we present and discuss the above-mentioned computational results, see Table LABEL:Table-chi-1 and Remark 5.3.
1. Varieties Isogenous to a Product
@@ 1.1. Group-Theoretic Description of an Unmixed VIP
$\overline{\psi_{i}}\colon G_{i}=G/K_{i}\hookrightarrow\operatorname{Aut}(C_{i})$
@@
is the set of group elements of $H$ with non-empty fixed locus on the curve $C$. It is hence called the stabilizer set of $V$.
@@
1.2. Generalities on Pluricanonical Maps of a VIP
In this section, we collect some general results on the $m$-canonical map of a VIP. Our main goal is to study its birationality for a given $m\geq 1$. Since such a variety is of general type, there exists a positive integer $m_{0}$ such that this map is birational for all $m\geq m_{0}$.
[…]
(1.2) $Z_{m}\cap V(x_{0},\dotso,x_{k})=\Phi_{m}^{Z}\left(V(s_{0},\dotso,s_{k})\right).$
Proposition 1.7.
Proof.
Note that $K_{X}^{\otimes m}$ being basepoint free amounts to saying that $V(s_{0},\dotso,s_{k})=\emptyset$. This is equivalent by (1.2) to $Z_{m}$ not meeting the centre of the projection pp. In other words, $\Phi_{m}^{X}$ is a morphism if and only if $p_{|Z_{m}}$ is a morphism. In this case $p_{|Z_{m}}$ is a finite morphism, therefore from diagram (1.1) it follows that $\Phi_{m}^{X}$ must be finite. ∎
Corollary 1.8.
Proposition 1.9.
  1. (1)
  2. (2)
    If one of the curves $C_{i}$ has genus $2$, then $\Phi_{2}^{X}$ is not birational.
Proof.
(1) Suppose w.l.o.g. that $C_{1}$ is hyperelliptic. Then the map $\Phi_{1}^{C_{1}}\colon C_{1}\to\mathbb{P}^{1}$ is a double cover, thus $Z_{1}\cong\mathbb{P}^{1}\times\overset{\sim}{C_{2}}\times\dotso\times\overset{\sim}{C_{n}}$. If $\Phi_{1}^{X}$ is birational, then from diagram (1.1) we get a dominant rational map
Proposition 1.10.
Proof.
Since $K_{S}$ is ample and $K^{2}_{S}=8\chi(\mathcal{O}_{S})\geq 8$, the result follows by [B-H-P-V, Theorem VII.5.1]. ∎
Remark 1.11.
By [CCZ07, Theorem 1.1] the $m$-canonical map of a minimal Gorenstein projective threefold of general type with canonical singularities is birational for $m\geq 5$. This holds in particular for threefolds isogenous to a product since they are smooth and their canonical bundle is ample.
Now we investigate the birationality of $\Phi_{4}^{X}$ for a threefold $X$ isogenous to a product. For this purpose, we recall that a $(1,2)$ surface $S$ is a nonsingular projective surface of general type whose minimal model $S^{\prime}$ has $K^{2}_{S^{\prime}}=1$ and $p_{g}(S^{\prime})=2$. By [Hor76II, Theorem 4.8] a $(1,2)$ surface is simply connected.
Theorem 1.12.
Proof.
Since $p_{g}(X)\geq 5$, [CH21, Theorem 1.1] implies that $\Phi_{4}^{X}$ is birational if and only if there is no smooth threefold $Y$ birational to $X$ admitting a fibration $Y\to B$, where $B$ is a smooth curve and the general fibre $S$ is a $(1,2)$-surface. By contradiction, assume that such a threefold $Y$ exists. Denote by $S^{\prime}$ the birational image in $X$ of the general fibre $S\subset Y$ and by $\hat{S}$ a resolution of singularities of $S^{\prime}$. Notice the $\hat{S}$ is simply connected since it is birational to the $(1,2)$ surface $S$. Hence, the map $\hat{S}\to S^{\prime}\subset X$ lifts to the universal cover of $X$, which is the product $\Delta^{3}$ of three unit disks:
[…]
@@ 2. Pluricanonical Representations of an Unmixed VIP
(2.1) ${C_{1}\times\dotso\times C_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{1}^{\prime}\times\dotso\times C_{n}^{\prime}}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$
@@
(2.2) $(g_{1},1,\dotso,1)\cdot(g,g,\dotso,g)\cdot(g_{1}^{-1},1,\dotso,1)\cdot(h^{-1},h^{-1},\dotso,h^{-1})\in K_{1}\times\dotso\times K_{n}.$
In other words, by the minimality of the realization we have that $gh^{-1}\in K_{2}\cap\dotso\cap K_{n}=\{1\}.$ Thus, $g=h$, which implies by (2.2) that $g_{1},g]\in K_{1}$. Since $g_{1},g\in G$ are arbitrary, this shows that $G,G]\leq K_{1}$. Similarly, one can show that $G,G]\leq K_{i}$, $i=2,\dotso,n$. Thus, we get that $G,G]=\bigcap_{i=1}^{n}K_{i}=\{1\}$, i.e. $G$ is abelian. ∎
@@
(2.3) $\rho\colon G_{1}\times\dotso\times G_{n}\to\operatorname{GL}\big(H^{0}(C_{1}\times\dotso\times C_{n},K_{C_{1}\times\dotso\times C_{n}}^{\otimes m})\big).$
@@
As for (1), let $v\in V^{H}$, $g\in N$ and $h\in H$. Then $g^{-1}hg\in H$, and therefore $\rho(g^{-1}hg)(v)=v$, which implies $\rho(h)\rho(g)(v)=\rho(g)(v)$, i.e. $\rho(g)(v)\in V^{H}$.
As for (2), by Maschke’s lemma there exists an $N$-invariant complement $W$ of $V^{H}$, namely $V=V^{H}\oplus W$. A constituent $n_{\chi}\chi$ of $\chi_{\rho_{|N}}$ belongs to the character of $V^{H}$ if and only if $\chi_{|H}$ is a multiple of the trivial character $\chi_{triv}$. ∎
By Künneth formula the action $\rho$ from (2.3) is the tensor product of the representations
@@
whose characters can be computed using the Chevalley-Weil formula (see [CheWeil34, Weil35]) in terms of the generating vectors $V_{i}$ of an algebraic datum of $X$. Thus, by the Chevalley-Weil formula and by Lemma 2.3 we recover from an algebraic datum of $X$ the character $\chi_{m}$ of the representation $\rho_{m}$.